# Causality-enhanced Decision-Making for Autonomous Mobile Robots in Dynamic Environments

### v3

@@ Abstract
The growing integration of robots in shared environments—such as warehouses, shopping centres, and hospitals—demands a deep understanding of the underlying dynamics and human behaviours, including how, when, and where individuals engage in various activities and interactions. This knowledge goes beyond simple correlation studies and requires a more comprehensive causal analysis. By leveraging causal inference to model cause-and-effect relationships, we can better anticipate critical environmental factors and enable autonomous robots to plan and execute tasks more effectively. To this end, we propose a novel causality-based decision-making framework that reasons over a learned causal model to predict battery usage and human obstructions, understanding how these factors could influence robot task execution. Such reasoning framework assists the robot in deciding when and how to complete a given task. To achieve this, we developed also PeopleFlow, a new Gazebo-based simulator designed to model context-sensitive human-robot spatial interactions in shared workspaces. PeopleFlow features realistic human and robot trajectories influenced by contextual factors such as time, environment layout, and robot state, and can simulate a large number of agents. While the simulator is general-purpose, in this paper we focus on a warehouse-like environment as a case study, where we conduct an extensive evaluation benchmarking our causal approach against a non-causal baseline. Our findings demonstrate the efficacy of the proposed solutions, highlighting how causal reasoning enables autonomous robots to operate more efficiently and safely in dynamic environments shared with humans.
@@ 1 Introduction
Autonomous mobile robots are a game changer for the progress and development of various sectors, including industry, logistics, agriculture, and healthcare. However, operating in dynamic environments and workspaces shared with people introduces several challenges. In particular, unexpected events and human encounters can negatively affect the capacity of a robot to fulfill its task safely and autonomously. Understanding the causal effect of the robot’s actions enables it to better assess and choose them, which is crucial for improving both the efficiency and safety of its operations. This motivates the need for intelligent solutions to reason about the causal relationships between the environment and human-robot interactions, improving robot’s efficiency and guaranteeing human safety. Unsurprisingly, the importance of cause-and-effect reasoning in AI and robotics has become strategically important for national and international research programmes111euRobotics, “A Unified Vision for European Robotics”, https://eu-robotics.net/strategy/, [^2] National Robotics Initiative 3.0, “Innovations in Integration of Robotics”, https://www.nsf.gov/pubs/2021/nsf21559/nsf21559.htm[1, 2].
Knowing the causal structure of a system provides a significant advantage in many robotics applications. Indeed, research in causal inference, which includes causal discovery and reasoning, has gained increasing attention in recent robotics literature [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. However, to the best of our knowledge, no prior work
@@
has explored causal inference for robot decision-making in the context of dynamic environments and human-aware navigation. The latter in particular typically relies on predictive models of human motion, which are often reactive and do not take into account robot and human goals, or how contextual factors can influence them. As a result, autonomous robot navigation in these environments can lead to potential safety and efficiency issues.
@@
  • a novel causal inference framework for robotic decision-making, integrated within the Robot Operating System (ROS), to improve task-execution efficiency and safety in dynamic human-shared environments;
  • a new Gazebo-based simulator, called PeopleFlow [^3] PeopleFlow GitHub repository: https://github.com/lcastri/PeopleFlow, which generates complex human and robot spatial behaviours considering various contextual factors (e.g., dynamic changes based on different goals and times of the day), as shown in Fig. 1;
  • a thorough experimental evaluation of the causal framework in a complex warehouse simulation to demonstrate its feasibility and potential for real-world robotic applications.
@@
The paper is structured as follows: basic concepts about causal discovery and related work in the robotics fields are presented in Section 2; Sections 3 and 4 explains the details of our causality-based decision making framework and PeopleFlow; Section 5 presents the application and results of our approach in a simulated warehouse-like environment; Section 6 justifies the choice of causality over standard Bayesian inference; finally, Section 7 concludes the paper discussing achievements and future work.
@@ 2 Related Work
Causal Discovery in Robotics:
@@
Causal discovery involves identifying the underlying causal structure from data, either through observational methods, interventional experiments, or a combination of both. It aims to uncover cause-and-effect relationships between variables rather than mere correlations. Numerous methods have been developed for causal discovery from both static and time-series data [16, 17]. However, for robotics applications, where time-series sensor data is common, methods tailored specifically for time-dependent causal discovery are essential. Among the various approaches available in the literature, the time-series variation of the Peter and Clark (PC) [18] algorithm, such as PC Momentary Conditional Independence (PCMCI) [19], has been widely used in many applications, including robotics [20, 21, 6]. PCMCI has recently been extended with several enhancements. PCMCI+ [22] enables the discovery of simultaneous dependencies. Filtered-PCMCI (F-PCMCI) [7] integrates a transfer entropy-based feature-selection module to improve causal discovery by focusing on relevant variables. Joint-PCMCI+ (J-PCMCI+[23] supports causal structure learning from multiple observational datasets using context variables. Latent-PCMCI (LPCMCI) [24] allows causal discovery in the presence of latent confounders. CAnDOIT [25], which stands for Causal Discovery with Observational and Interventional Data from Time Series, extends LPCMCI by integrating both observational and interventional data for improved causal discovery. Additionally, PCMCI and F-PCMCI have been integrated into a ROS-based causal discovery framework, ROS-Causal [26], facilitating their applicability in robotic systems.
Causal Reasoning in Robotics:
@@
Once the causal structure is established, causal reasoning enables inference, prediction, and decision-making by leveraging the discovered causal relationships to estimate the values of variables within the causal model, assess the effects of interventions, and evaluate counterfactual scenarios. Structural Causal Models (SCMs) have been employed to understand how humanoid robots interact with tools [4]. PCMCI and F-PCMCI have been used to establish causal models for underwater robots navigating toward target positions [5] and to predict human spatial interactions in social robotics [6, 7]. Moreover, causality-based approaches have been explored in various robotics domains, including robot imitation learning, manipulation and explainable HRI [8, 9, 10, 14, 15].
@@
Despite the growing use of causality in robotics, most existing approaches primarily focus on causal discovery and the prediction of time-series variables within the causal model. While some recent works [11, 12, 13] have explored causal reasoning for robot manipulation tasks, drone applications and planning, the use of causal models for decision-making in human-centered environments remains largely unexplored. In contrast, our work focuses on integrating causal reasoning into the robot’s decision-making process in dynamic environments shared with humans. By leveraging causal inference, our framework enables robots to make causality-aware decisions, enhancing both task execution efficiency and human safety in HRSI scenarios.
Human Robot Spatial Interaction:
@@
Several studies have highlighted the importance of considering the coexistence of humans and robots in shared environment, whether they are collaborating or performing individual tasks [27, 28, 29]. In this work, we focus on the spatial aspect of the interactions between human-human, human-robot and human-object [30]. Recently, various methods have been proposed to model the relations between human motion behaviours and spatial interactions [31, 32, 33]. However, these approaches have not explicitly considered causal relationships between spatial variables. Some recent studies [34, 6, 7] have developed causal models for modelling HRSI.
@@
Most existing datasets primarily focus on human motion trajectories rather than capturing how contextual elements shape human-robot interactions. For instance, THÖR [35] and its extension THÖR-MAGNI [36] provide motion trajectories recorded in controlled indoor environments using ceiling-mounted cameras, while the ATC Pedestrian Tracking dataset [37] tracks pedestrians in a shopping mall atrium using 3D range sensors. Similarly, the JackRabbot Dataset and Benchmark (JRDB) [38] captures human poses from a mobile robot’s ego-perspective using RGB cameras and 3D LiDAR. While these datasets offer valuable insights into human motion patterns, they do not include contextual variability, such as how different environmental settings, external agents, or task constraints influence interaction dynamics.
To address this limitation, simulation environments offer a practical solution for generating diverse, controlled, and repeatable HRSI scenarios. By systematically varying contextual factors, simulators enable the exploration of how different conditions influence human-robot interactions, effectively filling the gap left by existing datasets. SEAN 2.0 [39] provides a high-fidelity, Unity-based simulation environment tailored for pedestrian interactions, making it well-suited for training vision-based robotics algorithms in dynamic social settings. In contrast, MengeROS [40], CrowdNav [41], and SocialGym [42] focus on 2D grid-based visualisations, which, while useful for algorithmic benchmarking, may lack the realism needed to model complex human behaviours in rich environments. ROS-Causal_HRISim [^4] https://github.com/lcastri/ROS-Causal_HRISim [26] takes a different approach by enabling causal analysis in HRI scenarios, allowing both observational data collection and targeted interventions involving humans and robots.
Our new Gazebo-based simulator, PeopleFlow, builds upon ROS-Causal_HRISim to support context-sensitive human-robot spatial interactions. Unlike existing simulators, it does not only record HRSI data across diverse contexts but also models and dynamically assigns the roles of humans and robots based on probabilistic distributions conditioned on the context. This enhanced simulation framework offers the ability to record HRSI data across different contextual scenarios, in which people and robot have a role that depends on the context, enabling a more comprehensive analysis of how environmental conditions, agent behaviours, and task constraints influence human-robot interactions.
@@ 3 Causality-based Robot Decision Making Framework
Our approach aims to develop a causal inference engine [43] to enable causal reasoning in robotic decision-making. Specifically, we propose a framework that allows a robot to determine a path for completing a given task (e.g., delivering an object to a designated location) that represents a trade-off between factors relevant in its operating environment. These factors may include the current situation of the environment (e.g. people crowding an area at a specific time) or a robot-related constraint, such as battery level, that could influence task execution. Additionally, the framework should be capable of preemptively deciding whether to initiate or abort a task to avoid failures or inefficiencies.
@@ 3.1 Data Extraction Pipeline
Next, the data undergoes post-processing to extract relevant time-series information from the raw collected data, which is then stored for causal discovery analysis. To construct the causal inference engine, the extracted data must be compatible with inference mechanisms that often rely on discrete data representations. Therefore, we apply a discretisation step using pyAgrum [^5] https://pyagrum.readthedocs.io/en/1.17.2/’s built-in discretiser to transform continuous variables into discrete states while preserving essential statistical properties. Specifically, we use the quantile [44] and elbow method [45] approaches for discretisation. The quantile approach is particularly suitable when data is skewed or non-uniformly distributed, as in our case, whereas the elbow method helps to automatically determine the number of bins per variable.
@@ 3.2 Learning Pipeline
The learning phase comprises both causal discovery and parameters learning. It first focuses on uncovering the underlying causal relationships governing human and robot behaviours in the working environment. Once the causal structure is identified, the model’s parameters are estimated to enable reasoning over the retrieved causal relationships.
As previously explained, the causal discovery block takes as input the post-processed information from the data pipeline and performs the causal analysis to uncover the underlying causal structure governing the observed scenario. The analysis is conducted using CausalFlow [^6] https://github.com/lcastri/causalflow, a Python library that provides a collection of methods for causal discovery from time-series, as reviewed in Sec. 2. Embedding CausalFlow into our framework makes it adaptable to various scenarios. The choice of the causal discovery method depends on the dataset characteristics. For basic causal discovery, approaches such as PCMCI, DYNOTEARS, and VARLiNGAM can be applied. When the dataset requires feature selection, F-PCMCI is used, whereas L-PCMCI is more suitable when latent variables are present. If the dataset contains a mix of observational and interventional data, CAnDOIT is the appropriate method. In our case, since our dataset contains contextual variables, we use J-PCMCI+ [23].
Inspired by [46], the causal structure derived in the previous block is used together with the discretised data to learn the parameters of the causal inference engine. For this, we again employ pyAgrum5 for estimating the conditional probability distributions associated with each node in the causal graph, ensuring consistency with the causal structure identified in the discovery step. Unlike the approach in [46], which employs the Expectation-Maximization (EM) algorithm, we use Maximum Likelihood Estimation (MLE) [47], as our dataset does not contain missing data. These steps form the learning pipeline and are shown in Fig. 2.

### v4

@@ Abstract
The growing integration of robots in shared environments—such as warehouses, shopping centres, and hospitals—demands a deep understanding of the underlying dynamics and human behaviours, including how, when, and where individuals engage in various activities and interactions. This knowledge goes beyond simple correlation studies and requires a more comprehensive causal analysis. By leveraging causal inference to model cause-and-effect relationships, we can better anticipate critical environmental factors and enable autonomous robots to plan and execute tasks more effectively. To this end, we propose a novel causality-based decision-making framework that reasons over a learned causal model to assist the robot in deciding when and how to complete a given task. In the examined use case—i.e., a warehouse shared with people—we exploit the causal model to estimate battery usage and human obstructions as factors influencing the robot’s task execution. This reasoning framework supports the robot in making informed decisions about task timing and strategy. To achieve this, we developed also PeopleFlow, a new Gazebo-based simulator designed to model context-sensitive human-robot spatial interactions in shared workspaces. PeopleFlow features realistic human and robot trajectories influenced by contextual factors such as time, environment layout, and robot state, and can simulate a large number of agents While the simulator is general-purpose, in this paper we focus on a warehouse-like environment as a case study, where we conduct an extensive evaluation benchmarking our causal approach against a non-causal baseline. Our findings demonstrate the efficacy of the proposed solutions, highlighting how causal reasoning enables autonomous robots to operate more efficiently and safely in dynamic environments shared with humans.
@@ 1 Introduction
Autonomous mobile robots are a game changer for the progress and development of various sectors, including industry, logistics, agriculture, and healthcare. However, operating in dynamic environments and workspaces shared with people introduces several challenges. In particular, unexpected events and human encounters can negatively affect the capacity of a robot to fulfil its task safely and autonomously. Understanding the causal effect of the robot’s actions enables it to better assess and choose them, which is crucial for improving both the efficiency and safety of its operations. This motivates the need for intelligent solutions to reason about the causal relationships between the environment and human-robot interactions, improving robot’s efficiency and guaranteeing human safety. Unsurprisingly, the importance of cause-and-effect reasoning in AI and robotics has become strategically important for national and international research programmes111euRobotics, “A Unified Vision for European Robotics”, https://eu-robotics.net/strategy/, [^2] National Robotics Initiative 3.0, “Innovations in Integration of Robotics”, https://www.nsf.gov/pubs/2021/nsf21559/nsf21559.htm(euRobotics, 2024; National Science Foundation, 2021).
Knowing the causal structure of a system provides a significant advantage in many robotics applications. Indeed, research in causal inference, which includes causal discovery and reasoning, has gained increasing attention in recent robotics literature (Hellström, 2021; Brawer et al., 2020; Cao et al., 2021; Castri et al., 2022, 2023; Katz et al., 2018; Angelov et al., 2019; Lee et al., 2021; Cannizzaro et al., 2023b, a; Cannizzaro and Kunze, 2023; Love et al., 2024b, a). However, despite this growing attention, a notable gap persists between the potential benefits of modern causal inference and its actual application in real-world robotics. In fact, to the best of our knowledge, no prior work has explored causal inference for robot decision-making in the context of dynamic environments and human-aware navigation. The latter in particular typically relies on predictive models of human motion, which are often reactive and do not take into account robot and human goals, or how contextual factors can influence them. As a result, autonomous robot navigation in these environments can lead to potential safety and efficiency issues.
@@
  • a novel, end-to-end framework for robotic decision-making that, for the first time, integrates causal discovery from sensor data with causal reasoning for motion planning within the Robot Operating System (ROS) to improve task-execution efficiency and safety in dynamic human-shared environments;
  • a new Gazebo-based simulator, PeopleFlow [^3] PeopleFlow GitHub repository: https://github.com/lcastri/PeopleFlow, which generates complex spatial behaviours where agents adapt both reactively to dynamic changes and proactively by adjusting their goals based on contextual factors (e.g., time of day and scheduled tasks), as shown in Fig. 1;
  • a thorough experimental evaluation of the causal framework in a complex warehouse simulation to demonstrate its feasibility and potential for real-world robotic applications in comparison to a state-of-the-art non-causal baseline.
@@
The paper is structured as follows: basic concepts about causal discovery and related work in the robotics fields are presented in Section 2; Sections 3 and 4 explains the details of our causality-based decision making framework and PeopleFlow; Section 5 presents the application and results of our approach in a simulated warehouse-like environment; finally, Section 6 concludes the paper discussing achievements and future work.
@@ 2 Related Work
Causal Discovery in Robotics
@@
Causal discovery involves identifying the underlying causal structure from data, either through observational methods, interventional experiments, or a combination of both. It aims to uncover cause-and-effect relationships between variables rather than mere correlations. Numerous methods have been developed for causal discovery from both static and time-series data (Glymour et al., 2019; Assaad et al., 2022). However, for robotics applications, where time-series sensor data is common, methods tailored specifically for time-dependent causal discovery are essential. Among the various approaches available in the literature, the time-series variation of the Peter and Clark (PC) (Spirtes et al., 2000) algorithm, such as PC Momentary Conditional Independence (PCMCI) (Runge, 2018), has been widely used in many applications, including robotics (Runge et al., 2019; Saetia et al., 2021; Castri et al., 2022). PCMCI has recently been extended with several enhancements. PCMCI+ (Runge, 2020) enables the discovery of simultaneous dependencies. Filtered-PCMCI (F-PCMCI) (Castri et al., 2023) integrates a transfer entropy-based feature-selection module to improve causal discovery by focusing on relevant variables. Joint-PCMCI+ (J-PCMCI+(Günther et al., 2023) supports causal structure learning from multiple observational datasets using context variables. Latent-PCMCI (LPCMCI) (Gerhardus and Runge, 2020) allows causal discovery in the presence of latent confounders. CAnDOIT (Castri et al., 2024c), which stands for Causal Discovery with Observational and Interventional Data from Time Series, extends LPCMCI by integrating both observational and interventional data for improved causal discovery. Finally, PCMCI and F-PCMCI have been integrated into ROS-Causal (Castri et al., 2024a), a ROS-based framework that provides causal discovery only, facilitating their use in robotic systems but without supporting causal inference or reasoning.
Causal Reasoning in Robotics
@@
Once the causal structure is established, causal reasoning enables inference, prediction, and decision-making by leveraging the discovered causal relationships to estimate the values of variables within the causal model, assess the effects of interventions, and evaluate counterfactual scenarios. Structural Causal Models (SCMs) have been employed to understand how humanoid robots interact with tools (Brawer et al., 2020). PCMCI and F-PCMCI have been used to establish causal models for underwater robots navigating toward target positions (Cao et al., 2021) and to model and predict human spatial interactions in social robotics (Castri et al., 2022, 2023). The latter represent the first approaches to applying causal discovery for modelling HRSIs and reasoning in a forecasting task. However, they did not propose an end-to-end causal framework, nor were the predictions exploited for decision-making in human-aware robot navigation, as we do in this paper. Moreover, causality-based approaches have been explored in various robotics domains, including robot imitation learning, manipulation and explainable HRI (Katz et al., 2018; Angelov et al., 2019; Lee et al., 2021; Love et al., 2024b, a).
@@
Despite the growing use of causality in robotics, most existing approaches primarily focus on causal discovery and the prediction of time-series variables within the causal model. While some recent works (Cannizzaro et al., 2023b, a; Cannizzaro and Kunze, 2023) have explored causal reasoning for robot manipulation tasks and drone applications, the use of causal models for decision-making in human-centred environments remains largely unexplored. In contrast, our work focuses on integrating causal reasoning into the robot’s decision-making process in dynamic environments shared with humans. By leveraging causal inference, our framework enables robots to make causality-aware decisions, enhancing both task execution efficiency and human safety in HRSI scenarios.
Human Robot Spatial Interaction
@@
Several studies have highlighted the importance of considering the coexistence of humans and robots in shared environment, whether they are collaborating or performing individual tasks (Mukherjee et al., 2022; Vasconez et al., 2019; Jahanmahin et al., 2022). In this work, we focus on the spatial aspect of the interactions between human-human, human-robot and human-object (Dahiya et al., 2023). Recently, various methods have been proposed to model the relations between human motion behaviours and spatial interactions (Dondrup et al., 2015; Mghames et al., 2023a, b). However, these approaches have not explicitly considered causal relationships between spatial variables. Some recent studies (Liu et al., 2022; Castri et al., 2022, 2023) have developed causal models for modelling HRSI.
@@
Most existing datasets primarily focus on human motion trajectories rather than capturing how contextual elements shape human-robot interactions. For instance, THÖR (Rudenko et al., 2020) and its extension THÖR-MAGNI (Schreiter et al., 2024) provide motion trajectories recorded in controlled indoor environments using ceiling-mounted cameras, while the ATC Pedestrian Tracking dataset (Brščić et al., 2013) tracks pedestrians in a shopping mall atrium using 3D range sensors. Similarly, the JackRabbot Dataset and Benchmark (JRDB) (Martin-Martin et al., 2021) captures human poses from a mobile robot’s ego-perspective using RGB cameras and 3D LiDAR. While these datasets offer valuable insights into human motion patterns, they do not include contextual variability, such as how different environmental settings, external agents, or task constraints influence interaction dynamics.
To address this limitation, simulation environments offer a practical solution for generating diverse, controlled, and repeatable HRSI scenarios. By systematically varying contextual factors, simulators offer the opportunity to explore how different conditions may influence human-robot interactions, helping to address some of the limitations of existing datasets. SEAN 2.0 (Tsoi et al., 2022) provides a high-fidelity, Unity-based simulation environment tailored for pedestrian interactions, making it well-suited for training vision-based robotics algorithms in dynamic social settings. In contrast, MengeROS (Aroor et al., 2017), CrowdNav (Chen et al., 2019), and SocialGym (Holtz and Biswas, 2022) focus on 2D grid-based visualisations, which, while useful for algorithmic benchmarking, may lack the realism needed to model complex human behaviours in rich environments. ROS-Causal_HRISim [^4] https://github.com/lcastri/ROS-Causal_HRISim (Castri et al., 2024a) takes a different approach by enabling causal analysis in HRI scenarios, allowing both observational data collection and targeted interventions involving humans and robots.
Our new Gazebo-based simulator, PeopleFlow, extends our previous work ROS-Causal_HRISim to support context-sensitive human-robot spatial interactions. Unlike existing simulators, it does not only record HRSI data across diverse contexts but also models and dynamically assigns the roles of humans and robots based on probabilistic destination/goal station selections conditioned on context (e.g., time of the day, tasks to accomplish). This enhanced simulation framework offers the ability to record HRSI data across different contextual scenarios, in which people and robot have a role that depends on the context, enabling a more comprehensive analysis of how environmental conditions, agent behaviours, and task constraints influence human-robot interactions.
@@ 3 Causality-based Robot Decision Making Framework
Our approach introduces a novel, end-to-end framework for robotic decision-making, enabling an autonomous robot to reason about cause and effect for safer and more efficient navigation in dynamic environments. The core of this framework is a causal inference engine (Pearl and Mackenzie, 2018) that allows a robot to determine a path for completing a given task (e.g., delivering an object to a designated location), looking for a trade-off among relevant factors in its operating environment. These factors may include the current state of the environment (e.g. people crowding an area at a specific time) or a robot-related constraint, such as battery level, that could influence task execution. Additionally, the framework should be capable of preemptively deciding whether to initiate or abort a task to avoid failures or inefficiencies.
While prior work has used causal reasoning for time-series prediction (Castri et al., 2022, 2023), manipulation tasks (Cannizzaro et al., 2023b) and autonomous drone application (Cannizzaro et al., 2023a), our framework is the first to integrate these concepts into an end-to-end system for human-aware navigation. The primary contribution of this paper is this novel end-to-end causal architecture, which not only employs state-of-the-art algorithms to translate raw sensor data into a causally-based plan but is also architecturally designed to support long-term autonomy. This distinction is crucial: while a “two-stage” pipeline (offline causal discovery followed by a separate inference engine) might achieve similar initial performance, it is inherently static and requires human intervention for any single change. Our unified framework (shown in Fig. 2), being designed as a modular system, provides the foundation to overcome this dependency on human intervention, which is the key to enhancing long-term autonomy. For instance, if the environment evolves (e.g., new worker schedules or spatial layouts), a two-stage pipeline with a “frozen” causal inference engine leads to degraded performance until a human manually re-runs the entire discovery and deployment process. The proposed architecture could be easily extended to eventually manage its own model lifecycle by, for example, autonomously detecting when the engine is outdated and re-running the learning pipeline to adapt to the evolving environment.
@@ 3.1 Data Extraction Pipeline
Next, the data undergoes post-processing to extract relevant time-series information from the raw collected data, which is then stored for causal discovery analysis. To construct the causal inference engine, the extracted data must be compatible with inference mechanisms that often rely on discrete data representations. Therefore, we apply a discretisation step using pyAgrum [^5] https://pyagrum.readthedocs.io/en/1.17.2/’s built-in discretiser to transform continuous variables into discrete states while preserving essential statistical properties. Specifically, we use the quantile (Dimitrova et al., 2010) and elbow method (Dangeti, 2017) approaches for discretisation. The quantile approach is particularly suitable when data is skewed or non-uniformly distributed, as in our case, whereas the elbow method helps to automatically determine the number of bins per variable.
@@ 3.2 Learning Pipeline
The learning phase comprises both causal discovery and parameters learning. First, it first uncovers the underlying causal relationships governing human and robot behaviours in the working environment. This process assumes a predefined set of relevant variables based on domain knowledge. Once the causal structure is identified, the model’s parameters are estimated to enable reasoning over the retrieved causal relationships.
As previously explained, the causal discovery block takes as input the post-processed information from the data pipeline and performs the causal analysis to uncover the underlying causal structure governing the observed scenario. The analysis is conducted using CausalFlow [^6] https://github.com/lcastri/causalflow, a Python library that provides a collection of methods for causal discovery from time-series, as reviewed in Sec. 2. Embedding CausalFlow into our framework makes it adaptable to various scenarios. The choice of the causal discovery method depends on the dataset characteristics. For basic causal discovery, approaches such as PCMCI, DYNOTEARS, and VARLiNGAM can be applied. When the dataset requires feature selection, F-PCMCI is used, whereas L-PCMCI is more suitable when latent variables are present. If the dataset contains a mix of observational and interventional data, CAnDOIT is the appropriate method. In our case, since the dataset contains contextual variables, we use J-PCMCI+ (Günther et al., 2023). Given the data structure in our work—time-series with system variables and independent context variables—J-PCMCI+ is uniquely suited for this analysis. It is specifically designed to correctly model context variables as external factors that affect the system without being influenced by it. Applying other causal discovery methods from observational data, such as DYNOTEARS (Pamfil et al., 2020), VARLiNGAM (Hyvärinen et al., 2010), tsFCI (Entner and Hoyer, 2010), PCMCI (Runge, 2018), F-PCMCI (Castri et al., 2023), LPCMCI (Gerhardus and Runge, 2020), would incorrectly treat these external factors as standard system variables, thereby misrepresenting the underlying scenario. For this reason, we use only J-PCMCI+ in the following evaluation.
Inspired by (Raichev et al., 2024), the causal structure derived in the previous block is used together with the discretised data to learn the parameters of the causal inference engine. For this, we again employ pyAgrum5 for estimating the conditional probability distributions associated with each node in the causal graph, ensuring consistency with the causal structure identified in the discovery step. Unlike the approach in (Raichev et al., 2024), which employs the Expectation-Maximization (EM) algorithm, we use Maximum Likelihood Estimation (MLE) (Myung, 2003), as our dataset does not contain missing data. These steps form the learning pipeline and are shown in Fig. 2.